OpenStudy (anonymous):

i dont understand how to factor x^4 + 4x^3 - 17x^2 + 48x + 60 .... like... what are the steps...

OpenStudy (dumbcow):

--> (x^2 +ax + c)(x^2 +bx +d) cd = 60 ad+bc = 48 c+d = -17 a+b = 4

OpenStudy (anonymous):

are these cd.. ad+bc .. c+d.. and a+b just something i should remember when it comes to that factoring theory

OpenStudy (dumbcow):

not necessarily, it depends on the problem there is no guarantee it will factor either, in this case when you try to solve the system you may get no solution

OpenStudy (anonymous):

i see. i think i understand now. thank you

OpenStudy (dumbcow):

oh i messed up, it should be c+d+ab = -17 sorry

OpenStudy (anonymous):

factored out.. wouldnt it look like (x-1)(x+5)(x^2-12)

OpenStudy (dumbcow):

no not quite ... if you multiply that out you get x^4 + 4x^3 -17x^2 -48x +60

OpenStudy (anonymous):

thats what the equation is

OpenStudy (anonymous):

Another way to try this problem is to guess some of the roots using the Rational Roots theorem. If you get lucky you might find one, which results in being able to reduce the degree of your polynomial by 1.

OpenStudy (dumbcow):

good point ^^ you could also graph it to find any roots that could be factored out

OpenStudy (anonymous):

so if i am to solve this equation.. could i just leave the x values equal to + and - sqrt12

OpenStudy (anonymous):

and 1 and -5

OpenStudy (dumbcow):

wait it looks like you had a typo in your original question...you put "+48x" but it should be "-48x" otherwise you don't get a nice answer

OpenStudy (anonymous):

ahh . sorry about that.. . yes its -48x

OpenStudy (dumbcow):

in that case, you are correct sqrt12 can be simplified to 2sqrt3 though

OpenStudy (anonymous):

ok. ha. thanks again